OpenStudy (anonymous):

I need help with this problem: integrate sec^2(pix)tan^5(pix) dx

OpenStudy (anonymous):

\[\int\sec^2\pi x\tan^5\pi x~dx\] Let \(u=\tan\pi x\), so that \(du=\pi\sec^2\pi x~dx\), or \(\dfrac{1}{\pi}du=\sec^2\pi x~dx\): \[\frac{1}{\pi}\int u^5~du\]

OpenStudy (anonymous):

I first tried making u=tan(pix)

OpenStudy (anonymous):

ohhh, I messed up pretty early on. I didn't take out my constant

OpenStudy (anonymous):

Yep, gotta watch out for that chain rule.

OpenStudy (anonymous):

yeah, I tend to miss that :/

OpenStudy (anonymous):

then you just integrate u and sub tan(pix) back in

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

\[\frac{ 1 }{6 \pi }\tan ^{6}(pix)\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome!